Problem 4-4 (add the used formula)

PRESENT VALUE CALCULATIONS

What is the present value of the following future amounts?

Annual

Future Amount (SAR)

Interest

Years

900,000

5%

5

478,000

4%

12

279,000

12%

4

440,000

7%

15
Answer #1

Formula used: PV = FV/(1 + r)^n

Future Amount (SAR)

Annual Interest

Years

Present value

900,000

5%

5

PV = 900,000/(1 + 0.05)^5

PV = $705,173.549821613

478,000

4%

12

PV = 478,000/(1 + 0.04)^12

PV = $298,557.389699271

279,000

12%

4

PV = 279,000/(1 + 0.12)^4

PV = 298,557.389699271

440,000

7%

15

PV = 440,000/(1 + 0.07)^15

PV = $159,476.248642638

Can you please upvote? Thank You :-)